OpenStudy (anonymous):

Find the first 3 terms of the sequence that has a1 = 2, a2 = 6, and an = 3a(n − 1)a(n − 2).

OpenStudy (anonymous):

a3 = 3*a2*a1 = 36 a4 = 3*a3*a2 = 3*36*6 = 648 a5 = 3*a4*a3, etc